2. Conrad Jackson is a 28-year-old man who presents to the emergency department with severe fatigue and dehydration secondary to a 4-day history of vomiting. During the interview, he describes attending a family reunion and states that perhaps he ate something bad. Upon admission his vital signs are a temperature of 102.7°F, heart rate of 116 bpm, respiratory rate of 18 breaths/min, and blood pressure of 86/54 mm Hg. The nurse also notes the patient has dry mucous membranes and tenting of skin. The physician orders an IV to be started with 0.45% normal saline, and orders a serum electrolyte and an arterial blood gas. (Learning Objective 7) The following results are returned from the laboratory: Sodium (Na+) Potassium (K+) Chloride (CI) BUN Creatinine Glucose pH PaCO2 нсоз PaO2 O2 Sat 150 5.5 110 42 0.8 86 7.32 35 20 90 98%
a. What is your interpretation of this arterial blood gas sample? b. Explain the high potassium in this patient. c. Calculate the patients anion gap: d. What is the interpretation of this anion gap?
Answer #1

ANSWER

a) Patient's arterial blood gas report shows

                 * Hypernatremia (Increased sodium level)

                          Normal Na level = 135 -145 mEq/L.

                           Patient Na level = 150 mEq/L.

                * Hyperkalemia.(Increased pottasium level).

                         Normal pottasium level = 3.5 - 5 mEq/L.

                          Patient pottasium level = 5.5 mEq/L.

b) * Patient has high pottasium level due to dehydration which occur as

       a result of continous vomiting.

    * Dehydration result in changes in the concentration of electrolytes in the

      cells of the body and blood stream especially pottasium which leads to

    increased concentration of pottasium in the circulatory system.

c) Patient anion gap.

                Formula

                Anion gap (AG) = Na - (Cl + HCO3).

                      Na = 150 mEq/L.

                    Cl = 110 mEq/L.

                 HCO3 = 20 mEq/L.

   As per formula

                    AG = 150 - (110 + 20)

                           = 150 - 130

                          = 20 mEq /L

d) Interpretation of the anion gap.

              * Anion gap is a difference between measured cations

                and anions.

             * Normal range of anion gap is 3 to 11 mEq/L.

             * Patient anion gap is 20 mEq/L.

             * High anion gap value in this patient indicate intoxication

                or poisoning.